# Lindblad quantum dynamics from correlation functions of classical spin chains

### v1

@@ VI Conclusion
In summary, we have studied the Lindblad equation, as a central approach to the physics of open quantum systems. In this context, we have particularly focused on boundary-driven transport, where a steady state emerges in the long-time limit, which features a constant current and a characteristic density profile. The starting point of our study has been a recently suggested connection [72] between the dynamics in open and closed systems, which exits in certain cases. We have built on this connection for magnetization transport in the spin-1/2 XXZ chain with and without next-to-nearest-neighbor interactions, as integrability breaking perturbations. Specifically, we have shown that the time evolution of the open quantum system can be described by the use of classical correlation functions, as generated by the Hamiltonian equations of motion for classical real vectors. By comparing to exact numerical simulations of the Lindblad equation, we have demonstrated the accuracy of this description for a range of model parameters, but we have also pointed out its limitations by counterexamples. While this agreement is an interesting physical observation, it has also suggested that classical mechanics can be used to solve the Lindblad equation for comparatively large systems, which cannot be reached in a quantum mechanical treatment. Despite the approximate nature of this approach, one particular strength is the possibility to deal with small system-bath couplings, which are much harder to access on the basis of other numerical techniques. Furthermore, it seems to be possible that our approach can be applied to lattices in higher dimensions, as a promising direction of future research.

### v2

@@ VI Conclusion
In summary, we have studied the Lindblad equation, as a central approach to the physics of open quantum systems. In this context, we have particularly focused on boundary-driven transport, where a steady state emerges in the long-time limit, which features a constant current and a characteristic density profile. The starting point of our study has been a recently suggested connection [72] between the dynamics in open and closed systems, which exits in certain cases. We have built on this connection for magnetization transport in the spin-1/2 XXZ chain with and without next-to-nearest-neighbor interactions, as integrability breaking perturbations. Specifically, we have studied the question whether the time evolution of the open quantum system can be described by the use of classical correlation functions, as generated by the Hamiltonian equations of motion for classical real vectors. By comparing to exact numerical simulations of the Lindblad equation, we have found a good accuracy of such a description for a range of model parameters, but we have also pointed out counterexamples and its limitations for the quantitative extraction of transport coefficients. While this agreement is an interesting physical observation, it has also suggested that classical mechanics can be used to solve the Lindblad equation for comparatively large systems, which cannot be reached in a quantum mechanical treatment. Despite the approximate nature of this approach, one particular strength is the possibility to deal with small system-bath couplings, which are much harder to access on the basis of other numerical techniques.
An interesting question is the precise regime of validity of the quantum-classical correspondence of open systems within infinite-temperature spin-chain models. To answer it, one first has to know in which cases quantum and classical dynamics agree in the closed setup. Previous works like [64] suggest that such an agreement typically exists if the quantum model is nonintegrable, as it usually is for quasi-1D or 2D lattices. An agreement can also exist if the quantum model is integrable, as shown in Fig. 2. Yet, it necessarily requires that quasilocal conserved charges are of no relevance to the transport behavior [21], which is true for $\Delta>1$ in Fig. 2, but not for $\Delta<1$. An related point of view is that for $\Delta<1$ the system is more XY-like and quantum correlations become important.
Another interesting question is the existence of a quantum-classical correspondence of open systems outside infinite temperature and spin models. Since our present work relies on the relation in and around Eq. (15), one first has to answer if this relation can be generalized to other models and finite temperatures, which is a promising direction of future research. Apart from the generalization of Eq. (15), it is clear that quantum and classical dynamics cannot agree for all models and finite temperatures, neither in the closed nor in the open setup. Already for spin systems, the nature of low-energy excitations can be different in the quantum and classical case. Moreover, some systems have no classical counterpart.